OpenStudy (anonymous):

Can someone walk me through this? How do you find the length of segment AC on a number line? A is on -2 B is on 3 C is on 6

jimthompson5910 (jim_thompson5910):

AC = 6 - (-2) = 6+2 = 8 So the length of AC is 8 units

OpenStudy (anonymous):